OpenStudy (anonymous):

How do i do this math prolem 3x^4y^5 times 2xy^3

OpenStudy (e.mccormick):

Well, what does it ask for you to solve for? You said: \[3x^4y^5 \times 2xy^3\] That could be a simplify problem, or set to y, or set to x.

OpenStudy (anonymous):

this is a polynomial just multiply it but i dont know how lol

OpenStudy (anonymous):

well your going to have to get a simplfied answer

OpenStudy (e.mccormick):

Ah, exponent rules. You need to combind the exponents properly. For example \(x^2 times x^7 = x^{2+7}=x^9\)

OpenStudy (e.mccormick):

And the coeeficents, 2 and 3, do not forget to multiply them as well. Normal multiplication applies there.

OpenStudy (anonymous):

your answer should be 6x^5y^8

OpenStudy (anonymous):

Thank you!!!